# Expectations Versus Reality: Evaluating Intrusion Detection Systems in Practice

### v1

@@ I Introduction
There are several key challenges when standardising IDSs for comparison. The first one is using different datasets for testing. IDS solutions are often tested on diverse datasets, each with its unique characteristics and labelling methodologies. This variability complicates direct comparisons between different IDSs, as each system may be optimised for specific types of data. [3] Furthermore, the prevalent use of non-standardised datasets in testing makes it difficult to assess the generalisability and robustness of these systems. Where IDSs were tested on datasets constructed by the IDS developers, the performance seen in these tests may not generalise well when testing against other datasets. [4]
@@
The next significant challenge arises from the number of configuration options available in IDSs. These options, while allowing for customisation, can also lead to a lack of clarity in the evaluation methodology. Without a standardised approach to configuring these systems, results can vary widely, making it hard to determine the true effectiveness of an IDS under different scenarios. [5] One important factor in this particular area is that it remains unclear whether this is primarily due to the dataset, or the IDSs. Even for the most well-cited IDS datasets, there is literature available which questions the correctness of the data available. The intensive and complex process of dataset creation can often lead to error-prone output, presenting challenges in standardised testing if it does not also include extensive data wrangling and modification alongside it. [6]
@@
Another challenge is that IDSs commonly either take packets or flows. Where a dataset does not contain both of these formats, adapting it into the form expected by a given IDS is non-trivial, where the expected format is not the one provided by the dataset authors. This discrepancy presents challenges in obtaining satisfactory results when an IDS and dataset are incompatible without significant processing. [1] Our evaluation process was further complicated by the necessity of converting these datasets into formats compatible with various IDS solutions. This data wrangling could amplify the errors and inconsistencies inherent in the datasets. Such transformation processes can introduce additional noise and inaccuracies, thereby skewing the results of the IDS evaluations. For other important historical datasets that have been used to benchmark many IDSs (such as NSL-Cup 99), the pcap files may not be available at all. Where machine learning and autoencoder-based IDSs relied on particular features to make their assessment, there were difficulties where these files were unavailable or crucial features were not extracted, and utilising these systems became problematic. [1]
@@
Furthermore, the effectiveness of autoencoder-based IDSs, which rely on temporal and benign data to function optimally, is significantly hindered when pcap files are not categorised by scenario or if specific benign traffic data is absent. These systems require a baseline of ‘normal’ traffic against which to compare potentially malicious activity. Without this, training these models to accurately detect anomalies is challenging. The absence of scenario-specific or benign traffic data limits the ability of these systems to establish a normative profile of network activity, which is crucial for their anomaly detection mechanisms. In these cases, we were able to attempt to train the models on initial benign traffic in the dataset, but this often did not result in adequate performance and may not accurately represent a proper ‘baseline’ of traffic in the scenario. [7]
@@
Considering these issues, we propose a pipeline for the effective comparison of IDSs. Machine-learning based network IDSs (NIDS) have been increasing in popularity due to their ability to see more complex patterns, and thus display higher protection against unknown attacks than traditional IDSs. [7]. These benefits also come with an increase in complexity. In order to determine how flexible these systems are, we focus on NIDSs for the purpose of this study. To evaluate how well these NIDSs generalise, we use five NIDS datasets.
@@ II Related Work
When assessing NIDSs, there are two main components to be considered — how an IDS performs with a given dataset, and how datasets perform when run through a given NIDS. Concerning the latter component, work has been done in critically analysing the features of popular IDS datasets. Binbusayyis and Vaiyapuri [8] sought to find the optimal feature sets of 4 popular datasets using an ensemble method of statistical filters. Layeghy et al. [9] conducted an in-depth analysis of the statistical properties of benign traffic in the CICIDS2017, UNSW and TON-IOT datasets, and compared them with two real-world datasets. They found significant differences in the statistical features between the synthetic and real-world datasets and concluded that the evaluation of NIDS algorithms on synthetic datasets does not guarantee performance in real-world scenarios. Ghurab et al. [10] took an extensive look at the different NIDS datasets used for benchmarking, concluding that using recent datasets may be recommended due to their wider attack coverage, but note that all datasets may be appropriate in differing circumstances. However, since the datasets are not run through any IDSs in this study, there is no clear framework or pipeline provided that could be used to directly compare the performance of different IDSs on these datasets, which is crucial information for users who wish to adopt and use those IDSs in practice. Some other works have also looked at benchmarking IDSs, but many start at the evaluation steps with all common datasets and IDSs setup. [11, 12] Furthermore, some works propose frameworks for this benchmarking, but do not practically implement it. [13] Starting from this idealised view does not address the practical setup issues we encountered. Maseer et al. [14] investigated the performance of various machine learning algorithms, both supervised and unsupervised, on the CICIDS2017 dataset. This study evaluated the performance of various classical ML algorithms on the web-based attacks within CICIDS2017. Antunes et al. [15] evaluate the CSE-CIC-IDS2018 dataset and benchmark common deep learning methods, such as LSTM and CNN. However, these studies each consider a different single dataset in their benchmarking process, rather than a comprehensive evaluation across multiple datasets. As mentioned above, the characteristics of datasets differ, and it is suggested to use more recent datasets to observe results closer to the practical settings.
@@ VI-C Practical Implications for Deployment
    Need for Dynamic Testing Datasets: The disparities in IDS performance across datasets, such as HELAD’s high performance in a narrower dataset such as Stratosphere (F1 score of 0.9902) versus its lower performance in the more general UNSW-NB15 (F1 score of 0.0140), support calls for developing dynamic testing datasets that accurately simulate real-world conditions, including a mix of known threats and benign activities. [9]
@@
    Ongoing Adaptation and Update: The landscape of network threats is continually evolving, both datasets and IDSs can quickly fall behind current trends, this necessitates regular updates and adaptations of both IDS solutions and testing datasets to new threats and network conditions. [1]

### v3

@@ I Introduction
There are several key challenges when standardising IDSs for comparison. The first one is using different datasets for testing. IDS solutions are often tested on diverse datasets, each with its unique characteristics and labelling methodologies. This variability complicates direct comparisons between different IDSs, as each system may be optimised for specific types of data [3]. Furthermore, the prevalent use of non-standardised datasets in testing makes it difficult to assess the generalisability and robustness of these systems. Where IDSs were tested on datasets constructed by the IDS developers, the performance seen in these tests may not generalise well when testing against other datasets [4].
@@
The next significant challenge arises from the number of configuration options available in IDSs. These options, while allowing for customisation, can also lead to a lack of clarity in the evaluation methodology. Without a standardised approach to configuring these systems, results can vary widely, making it hard to determine the true effectiveness of an IDS under different scenarios [5]. One important factor in this particular area is that it remains unclear whether this is primarily due to the dataset, or the IDSs. Even for the most well-cited IDS datasets, there is literature available which questions the correctness of the data available. The intensive and complex process of dataset creation can often lead to error-prone output, presenting challenges in standardised testing if it does not also include extensive data wrangling and modification alongside it [6].
@@
Another challenge is that IDSs commonly either take packets or flows. Where a dataset does not contain both of these formats, adapting it into the form expected by a given IDS is non-trivial, where the expected format is not the one provided by the dataset authors. This discrepancy presents challenges in obtaining satisfactory results when an IDS and dataset are incompatible without significant processing [1]. Our evaluation process was further complicated by the necessity of converting these datasets into formats compatible with various IDS solutions. This data wrangling could amplify the errors and inconsistencies inherent in the datasets. Such transformation processes can introduce additional noise and inaccuracies, thereby skewing the results of the IDS evaluations. For other important historical datasets that have been used to benchmark many IDSs (such as NSL-Cup 99), the pcap files may not be available at all. Where machine learning and autoencoder-based IDSs relied on particular features to make their assessment, there were difficulties where these files were unavailable or crucial features were not extracted, and utilising these systems became problematic [1].
@@
Furthermore, the effectiveness of autoencoder-based IDSs, which rely on temporal and benign data to function optimally, is significantly hindered when pcap files are not categorised by scenario or if specific benign traffic data is absent. These systems require a baseline of ‘normal’ traffic against which to compare potentially malicious activity. Without this, training these models to accurately detect anomalies is challenging. The absence of scenario-specific or benign traffic data limits the ability of these systems to establish a normative profile of network activity, which is crucial for their anomaly detection mechanisms. In these cases, we were able to attempt to train the models on initial benign traffic in the dataset, but this often did not result in adequate performance and may not accurately represent a proper ‘baseline’ of traffic in the scenario [7].
@@
Considering these issues, we propose a pipeline for the effective comparison of IDSs. Machine-learning based network IDSs (NIDS) have been increasing in popularity due to their ability to see more complex patterns, and thus display higher protection against unknown attacks than traditional IDSs [7]. These benefits also come with an increase in complexity. In order to determine how flexible these systems are, we focus on NIDSs for the purpose of this study. To evaluate how well these NIDSs generalise, we use five NIDS datasets.
@@ II Related Work
When assessing NIDSs, there are two main components to be considered — how an IDS performs with a given dataset, and how datasets perform when run through a given NIDS. Concerning the latter component, work has been done in critically analysing the features of popular IDS datasets. Binbusayyis and Vaiyapuri [8] sought to find the optimal feature sets of 4 popular datasets using an ensemble method of statistical filters. Layeghy et al. [9] conducted an in-depth analysis of the statistical properties of benign traffic in the CICIDS2017, UNSW and TON-IOT datasets, and compared them with two real-world datasets. They found significant differences in the statistical features between the synthetic and real-world datasets and concluded that the evaluation of NIDS algorithms on synthetic datasets does not guarantee performance in real-world scenarios. Ghurab et al. [10] took an extensive look at the different NIDS datasets used for benchmarking, concluding that using recent datasets may be recommended due to their wider attack coverage, but note that all datasets may be appropriate in differing circumstances. However, since the datasets are not run through any IDSs in this study, there is no clear framework or pipeline provided that could be used to directly compare the performance of different IDSs on these datasets, which is crucial information for users who wish to adopt and use those IDSs in practice. Some other works have also looked at benchmarking IDSs, but many start at the evaluation steps with all common datasets and IDSs setup [11, 12]. Furthermore, some works propose frameworks for this benchmarking, but do not practically implement it [13]. Starting from this idealised view does not address the practical setup issues we encountered. Maseer et al. [14] investigated the performance of various machine learning algorithms, both supervised and unsupervised, on the CICIDS2017 dataset. This study evaluated the performance of various classical ML algorithms on the web-based attacks within CICIDS2017. Antunes et al. [15] evaluate the CSE-CIC-IDS2018 dataset and benchmark common deep learning methods, such as LSTM and CNN. However, these studies each consider a different single dataset in their benchmarking process, rather than a comprehensive evaluation across multiple datasets. As mentioned above, the characteristics of datasets differ, and it is suggested to use more recent datasets to observe results closer to the practical settings.
@@ VI-C Practical Implications for Deployment
    Need for Dynamic Testing Datasets: The disparities in IDS performance across datasets, such as HELAD’s high performance in a narrower dataset such as Stratosphere (F1 score of 0.9902) versus its lower performance in the more general UNSW-NB15 (F1 score of 0.0140), support calls for developing dynamic testing datasets that accurately simulate real-world conditions, including a mix of known threats and benign activities [9].
@@
    Ongoing Adaptation and Update: The landscape of network threats is continually evolving, both datasets and IDSs can quickly fall behind current trends, this necessitates regular updates and adaptations of both IDS solutions and testing datasets to new threats and network conditions [1].